# LionHeart: A Layer-based Mapping Framework for Heterogeneous Systems with Analog In-Memory Computing Tiles

### v1

@@ IV-F Temporal Drift and Evaluation Time Analysis
We study the impact of varying the evaluation time, $t_{eval}$, used during analog HWA retraining on the simulated inference time of the system. In Fig. 5, these metrics are reported for all configurations (except All Digital, which does not have a temporal dimension) using the ResNet8 network and CIFAR-100 dataset. For both sets of mappings, mostly unique intermediate layers are retained in analog. Due to the (relatively) small drop threshold, the first and last layers for each are in digital. It can be observed that the test accuracy decreases a function of the drift time. Moreover, as the proportion of analog MACs increases, the robustness to drift decreases. Most notably, for the All Analog configuration, independent of the evaluation time, the test accuracy decreases significantly (¿10%). For all other configurations, no significant degradation is observed until $t$¿1h. However, the temporal dependence of the test set accuracy is effected. Unsurprisingly, when a larger evaluation time is specified, the initial test set accuracy (at $t$=0) is also larger. While varied, for most configurations the robustness to drift is observed to improve as the evaluation time is increased. This is significant because, when our methodology is used, a hard constraint on the average behavior of the drop in accuracy can be enforced at a given drift (evaluation) time. Additionally, the test set accuracy for the FLMS configuration is consistently lower than for all LionHeart configurations.
V Conclusion
@@
In this paper, we presented LionHeart, an accuracy-driven framework for mapping DL workloads on hybrid analog-digital systems. The effectiveness of our method was demonstrated by deploying five different DNNs trained for image classification to heterogeneous PUs. Both the MAC ratio between analog and digital operations and the effect of temporal drift were considered to enforce varying user-specified constraints. Finally, the performance evaluation was carried out using a simulated ARMv8-based processor with tightly coupled AIMC tiles.
VI Acknowledgements
This work was supported by the European Commission under H2020 grant WiPLASH (GA 863337), the EC H2020 FVLLMONTI project (GA No. 101016776), and by the ACCESS – AI Chip Center for Emerging Smart Systems, sponsored by InnoHK funding, Hong Kong SAR. We also thank the IBM Research Analog AI team for their feedback and acknowledge computational support from AiMOS, an AI supercomputer made available by the IBM Research AI Hardware Center and Rensselaer Polytechnic Institute’s Center for Computational Innovations (CCI).

### v3

@@ VI Temporal Drift and Evaluation Time Analysis
We study the impact of varying the evaluation time, $t_{eval}$, used during analog HWA retraining on the simulated inference time of the system. In Fig. 9, these metrics are reported for all configurations (except All Digital, which does not have a temporal dimension) using the ResNet8 network and CIFAR-100 dataset. For both sets of mappings, mostly unique intermediate layers are retained in analog. Due to the (relatively) small drop threshold, the first and last layers for each are in digital. It can be seen that the test accuracy decreases as a function of the drift time. Moreover, as the proportion of analog MACs increases, the robustness to drift decreases. Most notably, for the All Analog configuration, independent of the evaluation time, the test accuracy decreases significantly ($>$10%). For all other configurations, no significant degradation is observed until $t>$1h. However, the temporal dependence of the test set accuracy is effected. Unsurprisingly, when a larger evaluation time is specified, the initial test set accuracy (at $t$=0) is also larger. While varied, for most configurations the robustness to drift is observed to improve as the evaluation time is increased. This is significant because, when our methodology is used, a hard constraint on the average behavior of the drop in accuracy can be enforced at a given drift (evaluation) time.
VII Comparison to Standard Hardware-Aware Retraining
Figure 10: Approximate Pareto front of the achieved LionHeart mappings, operational points of both variants of the Harmonica and FLMS methods, and the analog hardware-aware retraining ceiling for ResNet8 evaluated on CIFAR-10.
Finally, to compare LionHeart and other methods to standard hardware-aware retraining, we retrain all 1,024 mappings of ResNet8 and compute the exact Pareto front in Fig. 10, which is indicative of the ceiling or best case scenario when all analog layers are retrained at the same time using the same convergence criteria for all other experiments. It can be seen that for MAC ratio values smaller than 80%, LionHeart is extremely effective. For a MAC ratio of $\approx$60%, LionHeart exceeds this ceiling, indicating that progressively training analog layers can result in increased accuracy. This is in agreement with prior work [30].
VIII Conclusion
@@
In this paper, we have presented LionHeart, an accuracy-driven framework for mapping DL workloads on hybrid analog-digital systems. The effectiveness of our method was demonstrated by deploying five different DNNs trained for image classification to heterogeneous Processing Units. Both the MAC ratio between analog and digital operations and the effect of temporal drift were considered to enforce varying user-specified constraints. Performance evaluation was carried out for three DNNs using a simulated ARMv8-based processor with tightly coupled AIMC tiles. The optimal mappings of LionHeart were compared to related work, and it was demonstrated that LionHeart achieved state-of-the-art results. Finally, additional analysis was performed by investigating the effects of temporal drift and the variation of the evaluation time. Future work entails the hardware-based architecture-level exploration and investigation of the additional training time/complexity overhead that is incurred.
Acknowledgments
This work was supported in part by the the Swiss NSF Edge-Companions project (GA No. 10002812); in part by the EC H2020 FVLLMONTI Project under Grant 101016776; in part by the ACCESS—AI Chip Center for Emerging Smart Systems, sponsored by InnoHK funding, Hong Kong, SAR; in part by the Swiss State Secretariat for Education, Research, and Innovation (SERI) through the SwissChips Research Project; and in part by EC H2020 WiPLASH under Grant 863337.